OpenStudy (anonymous):

find the x-intercept of a parabola with vertex (-1,2) and y-intercept (0,-3) formula (y-k)=a(x-h)

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Well, the vertex of the parabola would be (h, k)

terenzreignz (terenzreignz):

So you can replace the h by -1 and the k by 2

OpenStudy (anonymous):

so -1-2=a(0-(-3)^2

terenzreignz (terenzreignz):

Yes, and solve for a.

terenzreignz (terenzreignz):

That doesn't end things yet, though, but do solve for a...

terenzreignz (terenzreignz):

WAIT

terenzreignz (terenzreignz):

Why did you replace y with -1?

terenzreignz (terenzreignz):

Let's start from the beginning...

OpenStudy (anonymous):

lol my bad

terenzreignz (terenzreignz):

\[\large (y-k) = a(x-h)^2\] Where (-1, 2) is the vertex, right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So, in general, In the equation\[\large (y-k) = a(x-h)^2\] The vertex is at the point (h, k) So you can replace (h,k) with (-1 , 2) And what do you get?

OpenStudy (anonymous):

0-(-1)=a(-3-2)^2

terenzreignz (terenzreignz):

Okay... that's the next step ahead.... now solve for a.

OpenStudy (anonymous):

1=a(25) im stuck here would i divide by 25?

terenzreignz (terenzreignz):

Yup. Don't worry about fractions....

OpenStudy (anonymous):

so a=1?

terenzreignz (terenzreignz):

No... if you divide both sides by 25, you get \[\large a =\frac1{25}\]

OpenStudy (anonymous):

so a is 1/25

terenzreignz (terenzreignz):

So, now you have your equation \[\huge y - 2 = \frac1{25}(x +2)^2\]

terenzreignz (terenzreignz):

Wait a sec... I seem to have gotten another mix-up

terenzreignz (terenzreignz):

Let's start from the beginning... again \[\large (y-k) = a(x-h)^2\]

OpenStudy (anonymous):

the answer has to be to the nearest tenth

terenzreignz (terenzreignz):

And don't skip steps... please :) for now, replace (h , k) with (-1 , 2)

OpenStudy (anonymous):

\[y-2=a(x-(-1)^2\]

terenzreignz (terenzreignz):

You got it backwards earlier \[\large y - 2 = a(x - (-1))^2\]

terenzreignz (terenzreignz):

Now, put in the y intercept, and replace (x , y) with (0 , -3)

OpenStudy (anonymous):

\[-3-2=a(0-(-1))\]

terenzreignz (terenzreignz):

Okay, don't forget the squared at the end... NOW solve for a.

OpenStudy (anonymous):

a=-5

terenzreignz (terenzreignz):

Yup... So now we have... \[\large y - 2= -5(x + 1)^2\] To get the x-intercepts, set y = 0, and solve for x.

OpenStudy (anonymous):

idk how to do that lol

terenzreignz (terenzreignz):

Set y = 0 \[\huge 0 - 2 = -5(x+1)^2\]

terenzreignz (terenzreignz):

\[\Large -2 = -5(x+1)^2\]

terenzreignz (terenzreignz):

Maybe you ought to try a bit harder... I'll be off for a bit, I'll be back in a few minutes..

OpenStudy (anonymous):

x=1

terenzreignz (terenzreignz):

@Mertsj maybe instead of expanding, you can divide both sides by -5 and just take the square roots?

OpenStudy (anonymous):

umm okay??

terenzreignz (terenzreignz):

Yeah... what that guy said ^ LOL divide both sides by -5 \[\huge \frac{-2}{-5}=(x+1)^2\]

terenzreignz (terenzreignz):

so, simplifies to... \[\large \frac25=(x+1)^2\]

terenzreignz (terenzreignz):

Taking the square roots of both sides gives you... \[\large \sqrt{\frac25}=x+1\]

terenzreignz (terenzreignz):

sorry \[\large \pm\sqrt{\frac25}=x+1\]

terenzreignz (terenzreignz):

And now, you can solve for x, quite simply :)

OpenStudy (anonymous):

.3675

terenzreignz (terenzreignz):

Don't forget the plus or minus...

OpenStudy (anonymous):

im so confused now

terenzreignz (terenzreignz):

Okay..., so solving for x gives us... \[x = -1 \pm \sqrt{\frac25}\] right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So, take both \[-1+\sqrt{\frac45}\]and\[-1-\sqrt{\frac45}\]

terenzreignz (terenzreignz):

And those are your x-intercepts.

OpenStudy (anonymous):

so (-5,0) (1,0)

OpenStudy (anonymous):

the answer has to be put like this (x1,y1) (x2,y2) to the nearest hundred

OpenStudy (anonymous):

??

terenzreignz (terenzreignz):

Yeah... the x intercepts look like these... (x1 , 0) and (x2 , 0 ) since the y-values on x-intercepts are always zero.